f(x) = x^2-9, g(x)=2, and h(x) = x-3
(f+g)(x)

Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39620)

Just substitute and simplify!

Answer by ikleyn(52803)

.

(f+g)(x) = x^2 - 9 + 2 = x^2 - 7.    ANSWER

Solved.